OpenStudy (anonymous):

How many 4-inch by 8-inch bricks are needed to build a walk 6 feet wide and 24 feet long?

OpenStudy (saifoo.khan):

576

OpenStudy (saifoo.khan):

20736/36

OpenStudy (anonymous):

A= 6*24 A=144ft^2 conver to in 144 * 12 = A = 1.728 in^2 A= 4*8 A= 32 in^2 1.728/32= 45 nedd 45 bricks

OpenStudy (anonymous):

click good answer ifI desever